OpenStudy (anonymous):

1. How do you write an equation of a line so that it is parallel or perpendicular to a given point and a given line?

OpenStudy (anonymous):

@Nurali

OpenStudy (paki):

y=mx+b parallel line equation

OpenStudy (owlcoffee):

Since we are not given any, let's take general things: Let's say we hace a line with equation y=(m1)x+b and a point T(k,n) wich is a point the parallel line passes through. By property of parallel lines, the slope of the first slope must be the same than the given. \[m_1 = m_2\] Now we know that the second equation goes through T(k.n). Then applying the equation: \[(y-y_1 ) = m(x-x_1)\] We know that m1 is equal to m2, and T(k.n) is a point, then: \[(y-n)=m_2 (x-k)\] Now, just simplify: \[y=m_2 x - m_2(x_1)-n\] must be the equation of the line parallel to the generalized line that was given, and the point also given. So basically I just deduced the slope with the property and transformed it to an equation using the given point and the equation that relates a slope and a point.

OpenStudy (anonymous):

You just identify the slope of the 1st line. Use that slope for parallel. Use the negative of that slope for perpendicular. Then, knowing slope and line, use point-slope form to build your second line. If I have the equation: \[2x=3y-9\] First I will solve it for y, to find m, my slope: \[3y = 2x + 9\\y = \frac{2x+9}{3}\\y = \frac{2}{3}x + 3\] So you know your slope, m, is 2/3. Now, you have another point that this 2nd line goes through. Let's say it's \[(1,1)\] Using point-slope form: \[(y-y_1) = m(x-x_1)\] \[y - 1 = \frac{2}{3}(x-1)\] Simplified: \[y = \frac{2}{3}x - \frac{2}{3} + 1\] \[y = \frac{2}{3}x + \frac{1}{3}\] This is the line that is parallel to the 1st, going through point (1,1).